# DualStreamFoveaNet: A Dual Stream Fusion Architecture with Anatomical Awareness for Robust Fovea Localization

### v4

@@ 1 Introduction
Detecting the location of the fovea is crucial in preventing irreversible damage to vision and is an essential part of automated fundus diagnosis [5, 6, 3]. However, several challenges arise in localizing fovea accurately. First, the surrounding retinal tissue’s color intensity makes the fovea’s dark appearance indistinguishable, and there are no anatomical landmarks in its vicinity [7, 8, 4]. Second, retinal diseases such as hemorrhages and microaneurysms can obscure the distinction between the fovea and retinal background, leading to inaccurate results [9, 10, 11, 12]. Similarly, bright lesions like exudates can change the fovea’s lightness to bright and lead to incorrect localization. Third, low light conditions and non-standard fovea locations during photography pose challenges in accurate fovea localization [13, 14, 15]. Blurred and poorly lit images make estimating macula difficult, and images with centrally located optic discs instead of macula can lead to opposite-side predictions compared to the ground truth. Given these challenges, a robust fovea localization method must take into account global image features to deliver accurate result.
@@
Fortunately, other anatomical structures outside the fovea, such as blood vessels, have proven to be useful for fovea localization in various studies [9, 10, 7, 8, 16, 17, 2, 3, 18, 19, 4, 20, 12, 15] and other lesion detection research [21, 22]. Previous works have employed carefully designed morphological methods to model the intricate anatomical relationships between the fovea and blood vessels [9, 2, 3, 4, 12]. However, these traditional methods based on morphological models may falter when faced with images exhibiting rare fovea positions and color intensity variations. While more recent studies have harnessed the power of deep learning to enhance performance, they typically rely solely on fundus images as input [23, 24, 11, 25, 13, 26, 14, 27] and are implemented on datasets containing few challenging images. These approaches result in three main shortcomings: 1) inadequate exploitation of the anatomical structure outside the macula, as only fundus images are used as input; 2) typical convolution-based architectures lacking the incorporation of global features; and 3) sensitivity to challenging cases, such as rare fovea positions and severe lesions.
@@
To tackle these challenges, we introduce a novel architecture, DualStreamFoveaNet (DSFN), for robust fovea localization. Drawing inspiration from TransFuser [28], our design features a dual-stream encoder that fuses multi-cue features and a decoder that generates result maps. By incorporating images from distinct cues (i.e., fundus and vessel distribution) as inputs for the encoder, we effectively utilize the anatomical structure outside the macula. In the encoder, we merge the multi-cue features of fundus and vessel images using transformer-based modules, dubbed the Bilateral Token Incorporation (BTI), which enables the modeling of global features and long-range connections for fovea localization, ensuring robust performance even in challenging images. Unlike TransFuser, which directly reduces and recovers token numbers by applying average pooling and bilinear interpolation methods—resulting in information loss—we employ TokenLearner and TokenFuser [29] within the BTI module. The attention mechanism of TokenLearner extracts self-learning spatial information from both cues, allowing our design to effectively exploit structural features along the optic disc and vessel distribution. Furthermore, the attention mechanism reduces the number of tokens in the BTI module, significantly cutting down on computational effort.
@@ 2.2 Deep Neural Networks in Fovea Localization
Deep learning has demonstrated superiority over traditional image processing and morphological techniques in many fields of medical image analysis, such as classification, segmentation and object localization [32, 33, 34, 35, 36]. Regarding the task of fovea localization, existing research can be broadly categorized into two primary domains: regression and segmentation.
@@ 3.1 Overall Architecture
The overall architecture of DSFN is depicted in Fig. 1. The encoder consists of two streams: the main stream and the satellite stream, which use ResNet34 and ResNet18 backbones respectively. The main stream extracts detailed features from fundus images, while the satellite stream extracts anatomical structure information from blood vessel distributions. Instead of using fundus images as input like the main stream, the satellite stream takes a vessel segmentation map generated by a pre-trained model. This pre-trained vessel segmentation model is built on the DRIVE dataset [45] using the TransUNet [42] architecture, which is identical to that used in [15].
@@
The dual-stream encoder, featuring four intermediate modules for multi-cue fusion, is inspired by PVT [46] and TransFuser [28]. Each stream’s backbone is divided into four convolutional blocks, comprising convolution and downsampling layers (Conv+Down). The resulting intermediate tensors ($F_{\mbox{\scriptsize{main}}}$ and $F_{\mbox{\scriptsize{satellite}}}$) are then fed into the BTI module, which includes a TokenLearner, $T$ consecutive Multi-Head Self-Attention (MHSA) layers, and a TokenFuser. The BTI module fuses multi-cue features and encodes long-range dependencies from both the fundus and vessel distribution. The output features are element-wise summed with skip-connected features and fed into the subsequent convolution and downsampling layers. Additionally, these output features from the BTI module are forwarded to RSU blocks for decoding operations.
@@
In contrast to the plain convolutional blocks commonly used in basic UNet decoders, the DSFN decoder employs four customized ReSidual U-blocks (RSU) [34] for feature fusion. Each block is a simple U-shaped structure for effective multi-scale feature incorporation. Qin et al. [34] demonstrate that the RSU block outperforms other embedded structures (e.g., plain convolution or residual-like, inception-like, and dense-like blocks) due to the enlarged receptive fields of the embedded U-shaped architecture. Furthermore, the superiority of the RSU structure as a decoder for incorporating multiple features has also been assessed by [15]. As illustrated in Fig. 1, RSU B4 serves as the bottleneck block between the encoder and decoder. The input to the other three RSU blocks is a concatenation of three types of features: (i) multi-scale skip-connected features from the main stream, (ii) multi-scale skip-connected features from the satellite stream, and (iii) the hidden feature decoded by the previous RSU block.
@@ 3.2.1 TokenLearner
The TokenLearner modules leverage spatial attention to adaptively generate tokens from the input feature maps for the subsequent MHSA layers. Let $\text{F}_{\text{in\_main}}\in\mathbb{R}^{h\times w\times c}$ and $\text{F}_{\text{in\_satellite}}\in\mathbb{R}^{h\times w\times c}$ be the input tensors from the two streams, where $h$, $w$, and $c$ represent the height, width, and channel of the corresponding BTI module. As shown in Fig. 2, the concatenated feature $\text{F}_{\text{in}}\in\mathbb{R}^{h\times w\times 2c}$ is first fed into the TokenLearner. We customize the TokenLearner [29] to contain a convolutional block containing two successive point-wise convolutional layers to reduce the dimensionality. This is followed by a flatten and a softmax operation to generate the spatial attention map $\text{F}_{\text{attn}}\in\mathbb{R}^{hw\times n}$, where $n$ is the number of learned tokens. Moreover, $\text{F}_{\text{in}}$ is processed by a point-wise convolutional layer, which is then flattened and transposed to generate a transformed feature map $\text{F}^{\prime}_{\text{in}}\in\mathbb{R}^{2c\times hw}$. The tokenization function is given by:
$\text{F}^{\prime}_{\text{in}}=\text{flatten}({h}(\text{F}_{\text{in}}))^{T},$ (1)
$\text{F}_{\text{attn}}=\text{softmax}(\text{flatten}({g}(\text{F}_{\text{in}}))),$ (2)
$\text{T}_{\text{L}}=\text{F}^{\prime}_{\text{in}}\text{F}_{\text{attn}},$ (3)
where ${h}(\cdot)$ represents a point-wise convolutional layer, ${g}(\cdot)$ is a block with two point-wise convolutional layers, and the learned tokens are denoted as $\text{T}_{\text{L}}\in\mathbb{R}^{2c\times n}$. Through its spatial attention mechanism, the TokenLearner module learns tokens that represent an informative combination of features from corresponding spatial locations across the input cues. In comparison to the 1024 tokens used in ViT and TransUNet [41, 42], we only retain $8\times 8$ tokens for each BTI module (given an input size of $512\times 512$). Since the computation of MHSA is quadratic to the number of tokens, the computational cost is significantly decreased. As demonstrated in our experiments, TokenLearner enables us to not only significantly reduce the number of tokens but also extract features related to the distributions of blood vessel for accurate fovea localization.
@@ 3.2.2 TokenFuser
The TokenFuser module reconstructs the tokens processed by the MHSA layers back to their original resolution, allowing the output feature ($\text{F}_{\text{out}}$) to be further processed by the remaining stages of the DSFN architecture. As shown in Fig. 2, it first transforms the MHSA output tokens to an intermediate token representation $\text{T}^{\prime}_{\text{L}}$ by a fully-connected (dense) layer. The TokenFuser then fuses this representation with the original input feature map $\text{F}_{\text{in}}$ in the following way:
$\text{F}^{\prime\prime}_{\text{in}}=\text{sigmoid}(r(\text{flatten}(\text{F}_{\text{in}})^{T}),$ (4)
$\text{F}_{\text{out}}=\text{reshape}((\text{T}^{\prime}_{\text{L}}\text{F}^{\prime\prime}_{\text{in}})^{T}),$ (5)
where tensors $\text{F}_{\text{in}},\text{F}_{\text{out}}\in\mathbb{R}^{h\times w\times 2c}$, $\text{F}^{\prime\prime}_{\text{in}}\in\mathbb{R}^{n\times hw}$ and $\text{T}^{\prime}_{\text{L}}\in\mathbb{R}^{2c\times n}$. ${r}(\cdot)$ represents a transformation involving two dense layers connected by an intermediate GeLU activation function. As we can see, the intermediate token representation $\text{T}^{\prime}_{\text{L}}$ is remapped to $\text{F}_{\text{out}}$ which has the same resolution as the original input feature map $\text{F}_{\text{in}}$. Next, $\text{F}_{\text{out}}$ is equally split to $\text{F}_{\text{out\_main}}$&$\text{F}_{\text{out\_satellite}}$, representing outputs corresponding to the main and satellite streams, respectively. These reconstructed features are then element-wise summed with the skip-connected features from each stream (i.e., $\text{F}_{\text{in\_main}}$&$\text{F}_{\text{in\_satellite}}$) to obtain the final outputs.
@@ 4.1 Datasets and Network Configurations
We also use a large-scale dataset (4103 images, named Tisu) which is collected from our cooperating hospital. All data have been desensitized with patients’ personal information removed. Compared to the Messidor and PALM datasets, Tisu is more challenging as it contains a larger number of fundus images with various abnormalities besides hemorrhages, microaneurysms, and exudates. The ground-truth fovea centers are determined by averaging the labels provided by three medical experts. The dataset is split into training and testing sets in a 4:1 ratio.

### v5

@@ 1 Introduction
Figure 1: Illustration of the shared global anatomical relationships among the fovea, optic disc, and main blood vessels, in normal (a),diseased (b, c), and poorly conditioned (d) retina images.
Detecting the location of the fovea is crucial in preventing irreversible damage to vision and is an essential part of automated fundus diagnosis [5, 6, 3]. Generally, the fovea is a small, central pit located in the macula region of the retina, responsible for sharp and detailed central vision. In retinal images, the fovea appears as a dark, circular area due to its high density of cone photoreceptors and lack of blood vessels, which enhances visual acuity. It is typically situated approximately 2.5 optic disc diameters temporal (toward the ear) from the center of the optic disc, aligned along the main axis of the retinal blood vessels (Fig. 1-a) [7, 2, 3].
@@
However, several challenges arise in localizing fovea accurately. First, the surrounding retinal tissue’s color intensity makes the fovea’s dark appearance indistinguishable, and there are no anatomical landmarks in its vicinity [8, 7, 4]. Second, retinal diseases such as hemorrhages and microaneurysms can obscure the distinction between the fovea and retinal background (Fig. 1-b and c), leading to inaccurate results [9, 10, 11, 12]. Similarly, bright lesions like exudates can change the fovea’s lightness to bright and lead to incorrect localization. Third, low light conditions and non-standard fovea locations during photography pose challenges in accurate fovea localization [13, 14, 15] (Fig. 1-d). Blurred and poorly lit images make estimating macula difficult, and images with centrally located optic discs instead of macula can lead to opposite-side predictions compared to the ground truth. Given these challenges, a robust fovea localization method must take into account global image features to deliver accurate result.
Fortunately, although the image conditions or retina diseases may reduce the distinctness between fovea and macula regions, the global anatomical structures outside the fovea, such as optic disc and blood vessels, have proved to be useful for fovea localization in various studies [9, 10, 8, 7, 16, 17, 2, 3, 18, 19, 4, 20, 12, 15] and other lesion detection research [21, 22]. The fovea center is typically situated along an imaginary line that extends from the center of the optic disc and runs through the macula. This line often aligns with the main superior and inferior temporal blood vessels, which branch out from the optic disc. Experienced clinicians utilize this spatial relationship between the optic disc and main vessels to approximate the fovea position when facing the aforementioned challenging retina images (Fig. 1-b, c and d) [9, 2, 23].
@@
Previous works have employed carefully designed morphological methods to model the intricate anatomical relationships between the fovea and blood vessels [9, 2, 3, 4, 12]. However, these traditional methods based on morphological models may falter when faced with images exhibiting rare fovea positions and color intensity variations. While more recent studies have harnessed the power of deep learning to enhance performance, they typically rely solely on fundus images as input [24, 25, 11, 26, 13, 27, 14, 28] and are implemented on datasets containing few challenging images. These approaches result in three main shortcomings: 1) inadequate exploitation of the anatomical structure outside the macula, as only fundus images are used as input; 2) typical convolution-based architectures lacking the incorporation of global features; and 3) sensitivity to challenging cases, such as rare fovea positions and severe lesions.
@@
To address these challenges, we introduce a novel architecture, DualStreamFoveaNet (DSFN), to adequately utilize the crucial global anatomical relationship (as shown in Fig. 1) for robust fovea localization. For many medical and biomedical research areas, the fusion of multi-branch features has proved to be a crucial technique for integrating information from different cues [29, 30, 31, 32, 33, 34]. Drawing inspiration from TransFuser [29], our design features a dual-stream encoder that fuses multi-cue features and a decoder that generates result maps. By incorporating images from distinct cues (i.e., fundus and vessel distribution) as inputs for the encoder, we effectively utilize the anatomical structure outside the macula. In the encoder, we merge the multi-cue features of fundus and vessel images using transformer-based modules, dubbed the Bilateral Token Incorporation (BTI), which enables the modeling of global features and long-range connections for fovea localization, ensuring robust performance even in challenging images. Unlike TransFuser, which directly reduces and recovers token numbers by applying average pooling and bilinear interpolation methods—resulting in information loss—we employ TokenLearner and TokenFuser [35] within the BTI module. The attention mechanism of TokenLearner extracts self-learning spatial information from both cues, allowing our design to effectively exploit structural features along the optic disc and vessel distribution. Furthermore, the attention mechanism reduces the number of tokens in the BTI module, significantly cutting down on computational effort.
@@ 2.2 Deep Neural Networks in Fovea Localization
Deep learning has demonstrated superiority over traditional image processing and morphological techniques in many fields of medical image analysis, such as classification, segmentation and object localization [38, 39, 40, 41, 42, 43]. Regarding the task of fovea localization, existing research can be broadly categorized into two primary domains: regression and segmentation.
@@ 3.1 Overall Architecture
The overall architecture of DSFN is depicted in Fig. 2. The encoder consists of two streams: the main stream and the satellite stream, which use ResNet34 and ResNet18 backbones, respectively. The main stream uses the larger ResNet34 to extract rich texture features from fundus images. In contrast, the satellite stream employs a simpler ResNet18 to extract key and sparse anatomical information from blood vessel distributions. Unlike the main stream that uses fundus images as input, the satellite stream takes a vessel segmentation map generated by a pre-trained model. This pre-trained vessel segmentation model is built on the DRIVE dataset [52] using the TransUNet [49] architecture, which is identical to that used in [15].
@@
The dual-stream encoder, featuring four intermediate modules for multi-cue fusion, is inspired by PVT [53] and TransFuser [29]. As shown in Fig. 2, each stream’s backbone is divided into four convolutional stages, comprising convolution, batch normalization, ReLU, and downsampling layers (Conv+Down). The resulting intermediate tensors ($\text{F}_{\mbox{\scriptsize{main}}}$ and $\text{F}_{\mbox{\scriptsize{satellite}}}$) are then fed into the BTI module, which includes a modified TokenLearner, $T$ consecutive Multi-Head Self-Attention (MHSA) layers, and a TokenFuser (Fig. 3).
Essentially, we have made four key improvements in the BTI module: 1) We innovatively utilize it to integrate multi-cue features generated from the convolution neural network. After rethinking the spatial mechanism [54, 55] used in TokenLearner, we find it well-suited for anatomical-based feature extraction. We adopt $8\times 8$ learned tokens instead of the original 8 tokens in TokenLearner before each ViT block, as the BTI module needs to adaptively learn features significantly related to anatomical structures (e.g., vessel distribution) for fovea localization. The $8\times 8$ tokens can retain sufficient spatial and other information for subsequent long-range dependency integration. 2) We progressively retain a greater number of channels (128, 256, 512, 1024) for the learned tokens in the deeper stages of the encoder’s BTI module. This approach is based on the observation that as the CNN training progresses, more abstract and high-level information is extracted from each stream. Retaining a greater number of channels allows for better integration and preservation of this information, enabling a gradual and effective fusion process. 3) We concatenate the input fundus and vessel features from the two streams first and then use two consecutive $1\times 1$ convolutional layers to remap the features along the channel. Since the non-vessel regions of concatenated vessel features are near zero, they explicitly provide clear clues for the spatial attention map to focus on anatomical structures. 4) After the TokenLearner, we innovatively use MHSA-based feature fusion to address the limitations of CNN-based feature fusion (i.e., limited receptive field and inductive bias of CNN). The MHSA block, through its self-attention mechanism, builds long-range dependencies, facilitating accurate fovea localization using global anatomical landmarks.
In the BTI module, anatomical-related features are adaptively learned, enabling efficient and effective fusion of multi-cue features from both fundus and vessel distributions by the subsequent multi-head self-attention mechanism. In other words, the adaptively learned features reduce the computational power required to encode long-range and global dependencies in the MHSA blocks, making it a practical choice for real-world applications.. The output features are element-wise summed with skip-connected features and fed into the next convolutional stage of each stream. Additionally, these output features from the BTI module are forwarded to RSU blocks for decoding operations.
@@
In contrast to the plain convolutional blocks commonly used in basic UNet decoders, the DSFN decoder employs four customized ReSidual U-blocks (RSU) [41] for feature fusion. Each block is a U-shaped structure for effective multi-scale feature incorporation. Qin et al. [41] demonstrate that the RSU block outperforms other embedded structures (e.g., plain convolution or residual-like, inception-like, and dense-like blocks) due to the enlarged receptive fields of the embedded U-shaped architecture. Furthermore, the superiority of the RSU structure as a decoder for incorporating multiple features has also been assessed by [15]. As illustrated in Fig. 2, RSU B4 serves as the bottleneck block between the encoder and decoder. The input to the other three RSU blocks is a concatenation of three types of features: (i) multi-scale skip-connected features from the main stream, (ii) multi-scale skip-connected features from the satellite stream, and (iii) the hidden feature decoded by the previous RSU block.
@@ 3.2.1 TokenLearner
The TokenLearner modules leverage spatial attention to adaptively generate tokens from the input feature maps for the subsequent MHSA layers. Let $\text{F}^{\text{main}}_{\text{in}}\in\mathbb{R}^{h\times w\times c}$ and $\text{F}^{\text{satellite}}_{\text{in}}\in\mathbb{R}^{h\times w\times c}$ be the input tensors from the two streams, where $h$, $w$, and $c$ represent the height, width, and channel of the corresponding BTI module. As shown in Fig. 3, the concatenated feature $\text{F}_{\text{in}}\in\mathbb{R}^{h\times w\times 2c}$ (Eq. 1) is first fed into the TokenLearner. We customize the TokenLearner [35] to contain a convolutional block containing two successive point-wise convolutional layers to reduce the dimensionality. This is followed by a flatten and a softmax operation to generate the spatial attention map $\text{F}_{\text{attn}}\in\mathbb{R}^{hw\times n}$ (Eq. 2), where $n$ is the number of learned tokens. Moreover, $\text{F}_{\text{in}}$ is processed by a point-wise convolutional layer, which is then flattened and transposed to generate a transformed feature map $\text{F}^{\prime}_{\text{in}}\in\mathbb{R}^{2c\times hw}$ (Eq. 3). Next, the learned tokens generated by spatial attention ($\text{F}_{\text{attn}}$) is given in Eq. 4.
${\text{F}_{\text{in}}=\text{Concat}(\text{F}^{\text{main}}_{\text{in}},\text{F}^{\text{satellite}}_{\text{in}})}$ (1)
$\text{F}_{\text{attn}}=\text{Softmax}(\text{flatten}({g}(\text{F}_{\text{in}})))$ (2)
$\text{F}^{\prime}_{\text{in}}=\text{Flatten}({h}(\text{F}_{\text{in}}))^{T}$ (3)
$\text{T}_{\text{L}}=\text{F}^{\prime}_{\text{in}}\text{F}_{\text{attn}}$ (4)
where ${g}(\cdot)$ is a block with two point-wise convolutional layers, ${h}(\cdot)$ represents a point-wise convolutional layer, and the learned tokens are denoted as $\text{T}_{\text{L}}\in\mathbb{R}^{2c\times n}$.
Through the spatial attention mechanism, the learned tokens represents an informative combination of features from corresponding spatial locations across the concatenated input cues. Given an input size of $512\times 512$, in comparison to the 1024 tokens used in ViT and TransUNet [48, 49], we only retain $8\times 8$ tokens for each BTI module. Since the computation of MHSA is quadratic to the number of tokens, the computational cost is significantly decreased. As demonstrated in our experiments, TokenLearner enables us to not only significantly reduce the number of tokens but also extract features related to the distributions of blood vessel for accurate fovea localization.
@@ 3.2.2 TokenFuser
The TokenFuser module reconstructs the tokens processed by the MHSA layers (i.e., $\mathbf{T}^{\prime}_{\text{L}}$ in Eq. 5) back to their original resolution, allowing the output feature ($\text{F}_{\text{out}}$) to be further processed by the remaining stages of the DSFN architecture. As shown in Fig. 3, the TokenFuser fuses $\mathbf{T}^{\prime}_{\text{L}}$ with the original input feature map ($\text{F}_{\text{in}}$) in the following way:
${\mathbf{T}^{\prime}_{\text{L}}=\text{MHSA}^{(t)}(\text{MHSA}^{(t-1)}(\cdots\text{MHSA}^{(1)}(\mathbf{T}_{\text{L}})\cdots)),t=12}$ (5)
$\text{F}^{\prime\prime}_{\text{in}}=\text{Sigmoid}(r(\text{flatten}(\text{F}_{\text{in}})^{T})$ (6)
$\text{F}_{\text{out}}=\text{Reshape}((\text{T}^{\prime}_{\text{L}}\text{F}^{\prime\prime}_{\text{in}})^{T})$ (7)
${\text{F}^{\prime\text{main}}_{\text{out}}=\text{F}_{\text{out}}[:,:,:c]\quad;\quad\text{F}^{\prime\text{satellite}}_{\text{out}}=\text{F}_{\text{out}}[:,:,c:]}$ (8)
${\text{F}^{\text{main}}_{\text{out}}=\text{F}^{\prime\text{main}}_{\text{out}}+\text{F}^{\text{main}}_{\text{in}}\quad;\quad\text{F}^{\text{satellite}}_{\text{out}}=\text{F}^{\prime\text{satellite}}_{\text{out}}+\text{F}^{\text{satellite}}_{\text{in}}}$ (9)
where tensors $\text{F}_{\text{in}},\text{F}_{\text{out}}\in\mathbb{R}^{h\times w\times 2c}$, $\text{F}^{\prime\prime}_{\text{in}}\in\mathbb{R}^{n\times hw}$ and $\text{T}^{\prime}_{\text{L}}\in\mathbb{R}^{2c\times n}$. ${r}(\cdot)$ represents a transformation involving two dense layers connected by an intermediate GeLU activation function. As we can see, the intermediate token representation $\text{T}^{\prime}_{\text{L}}$ is remapped to $\text{F}_{\text{out}}$ which has the same resolution as the original input feature map $\text{F}_{\text{in}}$ (Eq. 6 and Eq. 7). Next, $\text{F}_{\text{out}}$ is equally split to $\text{F}^{\prime\text{main}}_{\text{out}}$ and $\text{F}^{\prime\text{satellite}}_{\text{out}}$ (Eq. 8), representing outputs corresponding to the main and satellite streams, respectively. These reconstructed features are then element-wise summed with the skip-connected features from each stream (i.e., $\text{F}^{\text{main}}_{\text{in}}$ and $\text{F}^{\text{satellite}}_{\text{in}}$) to obtain the final output of each stream ($\text{F}^{\text{main}}_{\text{out}}$ and $\text{F}^{\text{satellite}}_{\text{out}}$ in Eq. 9).
@@ 4.1 Datasets and Network Configurations
Furthermore, we use a large-scale dataset (4103 images, named Tisu) which is a subset of data derived from a previously approved study, consists only of de-identified images to ensure subject privacy and data protection. Our current research accesses these images, maintaining confidentiality and ethical compliance. No identifiable subject information is included in the dataset used. Compared to the Messidor and PALM datasets, Tisu is more challenging as it contains a larger number of fundus images with various abnormalities besides hemorrhages, microaneurysms, and exudates. The ground-truth fovea centers are determined by averaging the labels provided by three medical experts. The dataset is split into training and testing sets in a 4:1 ratio.